OpenStudy (anonymous):

what are the steps to solving 2/5y- 6/7=1/4

OpenStudy (anonymous):

2/5y = (1/4 + 6/7) 140(2/5)y = 140[1/4 + 6/7) 56y =35+120 y=155/56=2.76786